Calculate the mass, in kg, of each sample.

1. 6.27×1026 cadmium atoms.

2. 1.93×1027 nickel atoms.

3. 6.43×1027 manganese atoms.

4. 1.89×1029 lithium atoms.

Calculate the number of atoms in each sample.

1. 11.781 g Cr

2. 31.750 g S

3. 23.286 g Pt

4. 82.515 g Sn

How many moles of atoms are in 7.00 g of 13C?
Answer #1

Calculate the mass, in kg, of each sample.

1. 6.27×1026 cadmium atoms.

Molar mass of Cd= 112.411 g/ mol

First calculate the number of moles

= 6.27*10^26 atoms * 1 mole/ 6.023*10^23atoms

= 1041.01 moles

Amount of 1041.01 moles Cd=

1041.01 moles *112.411 g/ mol

=117020.9 g

= 117.021 kg

2. 1.93×1027 nickel atoms.

Molar mass of Ni = 58.69 g/ mol

First calculate the number of moles

= 1.93*10^27 atoms * 1 mole/ 6.023*10^23atoms

= 3097.91 moles

Amount of 3097.91 moles Ni

3097.91 moles *58.69 g/ mol

=181816.5 g

= 181.82 kg

3. 6.43×1027 manganese atoms.

Molar mass of Mn= 54.94 g/ mol

First calculate the number of moles

= 6.43*10^27 atoms * 1 mole/ 6.023*10^23atoms

= 10675.7 moles

Amount of 10675.7 moles Mn

10675.7 moles *54.94 g/ mol

=586525.32 g

= 586.5 kg

4. 1.89×1029 lithium atoms.

Molar mass of Li= 6.941 g/ mol

First calculate the number of moles

= 1.89*10^29 atoms * 1 mole/ 6.023*10^23atoms

= 313797.11 moles

Amount of = 313797.11 moles Li

= 313797.11 moles* 6.941 g/ mol

=2178065.748 g

= 2178.066 kg